# LLM-Barber: Block-Aware Rebuilder for Sparsity Mask in One-Shot for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have grown significantly in scale, leading to a critical need for efficient model pruning techniques. Existing post-training pruning techniques primarily focus on measuring weight importance on converged dense models to determine salient weights to retain. However, they often overlook the changes in weight importance during the pruning process, which can lead to performance degradation in the pruned models. To address this issue, we present LLM-Barber (Block-Aware Rebuilder for Sparsity Mask in One-Shot), a novel one-shot pruning framework that rebuilds the sparsity mask of pruned models without any retraining or weight reconstruction. LLM-Barber incorporates block-aware error optimization across Self-Attention and MLP blocks, ensuring global performance optimization. Inspired by the recent discovery of prominent outliers in LLMs, LLM-Barber introduces an innovative pruning metric that identifies weight importance using weights multiplied by gradients. Our experiments show that LLM-Barber can efficiently prune models like LLaMA and OPT families with 7B to 13B parameters on a single A100 GPU in just 30 minutes, achieving state-of-the-art results in both perplexity and zero-shot performance across various language benchmarks. Code is available at https://github.com/YupengSu/LLM-Barber.
Introduction
Large language models (LLMs) have become a cornerstone in natural language processing (NLP) due to their impressive performance on various tasks. However, as these models increase in size and complexity, their deployment poses significant challenges due to extensive computational and storage demands. For instance, models such as GPT-175B (Brown et al. 2020), with 175 billion parameters, require vast resources, making it impractical for many applications. Therefore, efficient model compression strategies are crucial for deploying these powerful models in practical applications.
Model compression techniques commonly employ quantization and pruning to enhance model efficiency. Quantization reduces the precision of model parameters, while pruning removes less critical parameters. Traditional pruning methods, such as magnitude-based pruning (Han et al. 2015), directly trim weights based on their absolute values. While effective for smaller models, these methods often struggle with large-scale LLMs, resulting in suboptimal sparsity due to their inability to capture the complex interactions and importance of weights in such massive scalability. To address these limitations, recent post-training pruning techniques such as SparseGPT and Wanda have emerged.
Current pruning methods face two major challenges. First, as depicted in Figure 1(a), traditional layer-aware pruning methods focus on individual layers and neglect inter-layer dependencies within, leading to higher error accumulation (represented by blue arrows). In contrast, block-aware pruning, by considering groups of layers, captures inter-layer interactions to reduce error accumulation (represented by orange arrows). Second, as shown in Figure 1(b), conventional methods typically build the pruning mask once, ignoring the changes of weight significance in post-pruning stage. This oversight can lead to improper identification of salient weights, resulting in performance degradation.
@@ Introduction
To address these limitations, we propose LLM-Barber, a novel and straightforward approach designed to rebuild sparsity mask of pruned networks without requiring for retraining or weight reconstruction. Firstly, unlike layer-aware methods that are confined to local optimization and thus prone to significant error accumulation, as depicted by the blue arrows in Figure 1(a), LLM-Barber integrates pruning across both Self-Attention and MLP block. This approach mitigates error accumulation, as evidenced by the lighter orange arrows, facilitating global optimization and improved model performance. Secondly, LLM-Barber identifies weights that, although initially non-salient without a sparsity mask, gain significance in post-pruning. As shown in Figure 1(b), varying color shades represent the relative importance scores of different weights. LLM-Barber accurately identifies and rebuilds masks for high-score (deeply shaded) weights while pruning newly identified low-score (lightly shaded) weights. Thirdly, LLM-Barber utilizes a first-order Taylor series as pruning metric, which leverages weight-gradient multiplication for precise rebuilding of the sparsity mask, thereby enabling accurate pruning decisions and significantly reducing computational complexity compared to methods using second-order information. Finally, our one-shot approach surpasses iterative fine-tuning in efficiency while maintaining comparable accuracy. By rebuilding sparsity masks of pruned model at once, LLM-Barber offers a faster and more precise pruning solution. We validate the effectiveness of LLM-Barber through extensive experiments on the LLaMA and OPT families. Our method consistently outperforms existing post-training pruning techniques in both perplexity and zero-shot performance, establishing LLM-Barber as a state-of-the-art approach in this domain. To sum up, the key contributions are fourfold:
@@
  • Innovative Pruning Metric: We propose an innovative pruning metric based on the product of weights and gradients, leveraging first-order Taylor series for importance evaluation to reduce computational complexity compared to second-order Hessian-based approaches.
@@
Related Work
LLMs Pruning and Sparsity
Network pruning reduces deep neural networks by removing unnecessary weights. For LLMs, pruning methods are divided into parameter efficient fine-tuning and post-training approaches. Parameter Efficient Fine-tuning (PEFT) begins with a initialized sparse network and refines it through iterative processes (Liu et al. 2019). LoRA (Hu et al. 2021) adapts pre-trained models to specific tasks or domains by injecting trainable rank decomposition matrices. Dynamic Spase No Training (Zhang et al. 2024) minimizes the reconstruction error by iteratively pruning and growing weights. However, fine-tuning requires ample data and often leads to performance decline. Recent studies advance toward one-shot post-training pruning, showing substantial improvements. Post-training pruning removes weights from a pre-trained model. SparseGPT (Frantar and Alistarh 2023) uses Hessian-based metrics and subsequent residual weight updates, while Wanda (Sun et al. 2023) introduces a first-order pruning metric using weight-activation products. LLM-Barber employs a block-aware reconstruction approach and rebuilds masks with a novel pruning metric.
Model Compression Strategy
Compression is key to reducing the memory and computational demands of model. This work highlights the block-aware compression strategy over traditional layer-aware approaches. Layer-aware compression strategy began with Optimal Brain Damage (LeCun, Denker, and Solla 1989) and Optimal Brain Surgeon (Hassibi, Stork, and Wolff 1993). Building on these foundations, recent works like GPTQ (Frantar et al. 2023) further enhance layer-aware quantization using second-order information. Block-aware compression strategy generally offer better accuracy recovery in pruned models compared to layer-aware methods. For instance, APTQ (Guan et al. 2024) applies global quantization to attention mechanisms, enhancing model robustness, while BESA (Xu et al. 2024) uses block-wise sparsity allocation. Our method leverages block-aware pruning to optimize global performance across Self-Attention and MLP blocks, effectively balancing efficiency and accuracy.
Block-Aware Rebuilder for Sparsity Mask
Figure 2: The workflow of LLM-Barber. (a) illustrates the process of block-aware reconstruction error calculation and gradient back-propagation for each linear weight. (b) shows pruning metric computation and procedure for rebuilding sparsity mask.
Preliminaries
LLM pruning removes weights from dense networks to minimize output discrepancies, which is computationally intensive cross large-scale models, leading to address a layer-aware reconstruction problem (Hassibi, Stork, and Wolff 1993). This section reviews and reanalyses layer-aware reconstruction error and Taylor expansion at dense networks.
@@
$\mathbf{E}(\mathbf{\widehat{W}},\mathbf{X})=||\mathbf{WX}-\mathbf{\widehat{W}X}||_{2}^{2},$ (1)
@@
$\mathbf{E}(\mathbf{M},\mathbf{X})=||\mathbf{WX}-\mathbf{(W\odot M)X}||_{2}^{2}.$ (2)
The objective is to search for an optimal sparsity mask $\mathbf{M}$ where the pruning process reduces model complexity while preserving its predictive accuracy.
@@
$\mathbf{E}(\mathbf{\widehat{W}},\mathbf{X})=\mathbf{E}(\mathbf{\widehat{W}}_{\text{dense}},\mathbf{X})+\frac{\partial\mathbf{E}}{\partial\mathbf{W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (3)
Without a sparsity mask in dense model, we can simply assign all-one matrix to the mask M, thereby yielding the zeroth-order terms $\mathbf{E}(\mathbf{\widehat{W}_{\text{dense}}},\mathbf{X})=0$. The first-order derivative ${\partial\mathbf{E}}/\partial\mathbf{W}$ vanishes when training converged (Frantar and Alistarh 2023; Hassibi, Stork, and Wolff 1993), leaving only the computationally expensive second-order terms involving a large Hessian matrix which are challenging for layer-wise reconstruction and channel-wise independence assumption.
Block-Aware Rebuilder for Sparsity Mask
In this work, we depart from existing post-training pruning methods in three key aspects: Firstly, to address the layer-aware reconstruction problem that leads to exponentially accumulating biases, we adopt a block-aware reconstruction error and apply a divide-and-conquer strategy to mitigate errors and computational costs. Secondly, inspired by the pruning-and-growing111“Pruning” and “Growing” will both be used in the following text to refer to rebuild for sparsity masks. Pruning changes mask values from 1 to 0, whereas growing changes them from 0 to 1. operation (Mocanu et al. 2018; Zhang et al. 2024), we address the limitations of mask selection of pruning in dense networks due to the changeable significance of weights, by re-evaluating weight importance score in sparse networks and rebuilding the sparsity mask through targeted growth of salient weights and pruning of non-salient weights. Thirdly, our analysis reveals that with the advancement of LLMs, mask selection becomes increasingly critical in weight reconstruction. Thus, we prioritize the rebuilding of the sparsity mask and strip the reconstruction of weights. Base on these insights, we propose LLM-Barber, a Block-Aware Rebuilder for Sparsity Mask in One-Shot without any need for fine-tuning or retraining.
@@
$\mathbf{E(M,X)}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (4)
Evaluating reconstruction error across blocks, denoted as $\mathbf{\text{Block}(\cdot)}$, allows us to achieve a globally optimal solution in Self-Attention and MLP blocks rather than layer-wise. This new block-aware reconstruction formulation offers a more effective and cohesive strategy for selecting sparsity masks.
@@
$\mathbf{E(\widehat{W},X)}=\mathbf{E(\widehat{W}_{\text{sparse}},X)}+\mathbf{\frac{\partial E}{\partial W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (5)
The zeroth-order term of the Taylor expansion represents the reconstruction error after mask initialization:
$\mathbf{E(\widehat{W}_{\text{sparse}},X)}\!=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M_{i},X)}||_{2}^{2}.$ (6)
@@
where ||$\cdot$|| ⋅ | represents the absolute value operator, and $\mathbf{E}$ is denoted as the block-aware reconstruction error:
$\mathbf{E(M,X)}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (9)
@@
Pruning Granularity. Choosing the right pruning granularity is crucial (Sun et al. 2023). Traditional methods operate on a layer-wise (Han et al. 2015), input-wise (Frantar and Alistarh 2023), or output-wise (Sun et al. 2023) basis, which mainly address the layer-aware reconstruction problem, known for its output channel independence. Wanda’s output-wise ranking yields superior results compared to other methods. However, in LLM-Barber’s block-aware framework, output channel independence is no longer applicable. Thus, LLM-Barber extends consideration to block-wise granularity, prioritizing all linear layers within a block. Our analysis of the four distinct granularity levels shows that optimal granularity depends on the specific sparse mask initialization, with detailed results discussed in the Ablation StudyAblation Study.
Mask Rebuilding. With the block-aware reconstruction error $\mathbf{E}$, gradient information, sparsity metric, and granularity established, we proceed to rebuild the sparsity mask for each layer. Consider a cluster of weights $\mathbf{W^{c}}$ under a specific sparsity granularity and its corresponding sparsity mask $\mathbf{M^{c}}$. We define the growing criterion and the pruning criterion:
$gi,gj=\text{argmax}\ |\mathbf{W^{c}}|\!\cdot\!|(\partial\mathbf{E}/\partial\mathbf{W^{c}})|,\text{if}\ \mathbf{M^{c}}_{gi,gj}=0,$ (10)
$pi,pj=\text{argmin}\ |\mathbf{W^{c}}|\!\cdot\!|(\partial\mathbf{E}/\partial\mathbf{W^{c}})|,\text{if}\ \mathbf{M^{c}}_{pi,pj}=1.$ (11)
@@
Our experiments show that LLM-Barber identifies varying proportions of salient weights depending on the sparse mask’s initialization method. To control the extent of mask rebuilding, we introduce a hyperparameter $\alpha$, called the mask rebuilding ratio. The number of mask rebuilding pairs $N$ is calculated as:

### v2

# Graziano Chesi2, Ngai Wong2, Hao Yu1

@@ Abstract
Large language models (LLMs) have seen substantial growth, necessitating efficient model pruning techniques. Existing post-training pruning methods primarily measure weight importance in converged dense models, often overlooking changes in weight significance during the pruning process, leading to performance degradation. To address this issue, we present LLM-Barber (Block-Aware Rebuilder for Sparsity Mask in One-Shot), a novel one-shot pruning framework that rebuilds the sparsity mask of pruned models without any retraining or weight reconstruction. LLM-Barber incorporates block-aware error optimization across Self-Attention and MLP blocks, facilitating global performance optimization. We are the first to employ the product of weights and gradients as a pruning metric in the context of LLM post-training pruning. This enables accurate identification of weight importance in massive models and significantly reduces computational complexity compared to methods using second-order information. Our experiments show that LLM-Barber efficiently prunes models from LLaMA and OPT families (7B to 13B) on a single A100 GPU in just 30 minutes, achieving state-of-the-art results in both perplexity and zero-shot performance across various language benchmarks.
Index Terms:
Large Language Models, Post-training Pruning, Sparsity Mask Rebuilding, Low-resource Deployment
I Introduction
LLMs have become foundational in artificial intelligence due to their impressive performance on various tasks. However, the increasing size and complexity of models, such as GPT-175B [1] with 175 billion parameters, pose significant challenges related to extensive computational and storage demands. Consequently, efficient model compression strategies are crucial for enabling the practical deployment of these powerful models.
Current pruning methods face two major challenges. First, traditional layer-aware pruning methods focus on individual layers, neglecting inter-layer dependencies, which increases error accumulation (blue arrows in Figure 1(a)). In contrast, block-aware pruning considers groups of layers, effectively reducing this error (orange arrows). Second, as shown in Figure 1(b), conventional methods typically build the pruning mask only once, ignoring the changes of weight significance in post-pruning stage. This oversight can lead to improper identification of salient weights and subsequent performance degradation.
@@ I Introduction
To address these limitations, we propose LLM-Barber, a novel and straightforward approach rebuild sparsity mask of pruned networks without requiring for retraining or weight reconstruction.
Figure 2: The workflow of LLM-Barber. (a) illustrates the process of block-aware reconstruction error and gradient calculation for each linear weight. (b) shows pruning metric computation and sparsity mask rebuilding.
Firstly, unlike layer-aware methods that prioritize local optimization and are prone to error accumulation (illustrated by the blue arrows in Figure 1(a)), LLM-Barber integrates pruning across Self-Attention and MLP blocks. This approach mitigates error accumulation, as evidenced by the lighter orange arrows, promoting more global optimization. Secondly, LLM-Barber identifies weights that, although initially non-salient without a sparsity mask, gain significance in post-pruning. As shown in Figure 1(b), varying color shades represent the relative importance scores of different weights. LLM-Barber accurately rebuilds masks for high-score weights (deeply shaded) while pruning newly identified low-score weights (lightly shaded). thirdly, LLM-Barber employs the product of weights and gradients as pruning metric. While this metric has been applied in other contexts [2, 3], we are the first to leverage it in the context of post-training compression for LLMs. This enables precise rebuilding of the sparsity mask and significantly reduces computational complexity compared to methods [4] using second-order information. Finally, our one-shot approach efficiently rebuilds sparsity masks for pruned models, achieving comparable accuracy to iterative fine-tuning.
LLM-Barber demonstrates its effectiveness through experiments on the LLaMA and OPT families, consistently outperforms existing post-training techniques in both perplexity and zero-shot tasks. To sum up, the key contributions are fourfold:
@@
  • Innovative Pruning Metric: We are the first to apply the product of weights and gradients as a pruning metric within the context of LLM post-training pruning, leveraging first-order Taylor series for importance evaluation to reduce computational complexity compared to second-order Hessian-based approaches.
@@
II Related Work
II-A LLMs Pruning and Sparsity
Network pruning reduces deep neural networks by eliminating unnecessary weights, with methods categorized into parameter-efficient fine-tuning (PEFT) and post-training approaches. PEFT begins with an initialized sparse network and refines it through iterative processes [5]. LoRA [6] adapts pre-trained models to specific tasks by injecting trainable rank decomposition matrices. Dynamic Spase No Training [7] prunes and grows weights to minimize reconstruction error. However, fine-tuning often requires ample data and can degrade performance. Post-training, on the other hand, removes weights from a pre-trained model. SparseGPT [4] uses Hessian-based metrics and subsequent residual weight updates, Wanda [8] introduces a first-order metric using weight-activation products. LLM-Barber employs a block-aware reconstruction strategy and rebuilds masks with a novel pruning metric.
II-B Model Compression Strategy
Compression is key to reducing the memory and computational demands of model. Layer-aware strategies, originating from Optimal Brain Damage [9] and Optimal Brain Surgeon [10], have been enhanced by recent works like GPTQ [11] using second-order information. However, block-aware strategies typically yield better accuracy recovery. For instance, APTQ [12] applies global quantization to attention mechanisms, BESA [13] uses block-wise sparsity allocation. Our method leverages block-aware pruning to optimize global performance across blocks, effectively balancing efficiency and accuracy.
III LLM-Barber
III-A Preliminaries
LLM pruning removes weights from dense networks to minimize output discrepancies, which is computationally intensive cross large-scale models, leading to address a layer-aware reconstruction problem [10]. This section reviews and reanalyses layer-aware reconstruction error and Taylor expansion at dense networks.
@@ III-A Preliminaries
$\mathbf{E}=||\mathbf{WX}-\mathbf{\widehat{W}X}||_{2}^{2},$ (1)
@@
$\mathbf{E}=||\mathbf{WX}-\mathbf{(W\odot M)X}||_{2}^{2}.$ (2)
The goal is to find an optimal sparsity mask $\mathbf{M}$ that reduces model complexity while preserving predictive accuracy.
@@
$\mathbf{E}=\mathbf{E_{d}}+\frac{\partial\mathbf{E}}{\partial\mathbf{W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (3)
Without a sparsity mask in dense model, we can simply assign all-one matrix to the mask M, thereby yielding the zeroth-order terms $\mathbf{E_{d}}=0$. The first-order derivative ${\partial\mathbf{E}}/\partial\mathbf{W}$ vanishes when training converged [4, 10], leaving only the computationally expensive second-order terms involving a large Hessian matrix which are challenging for layer-wise reconstruction and channel-wise independence assumption.
III-B Block-Aware Rebuilder for Sparsity Mask
In this work, we depart from existing post-training pruning methods in three key aspects: Firstly, to address the layer-aware reconstruction problem that leads to exponentially accumulating biases, we adopt a block-aware reconstruction error and apply a divide-and-conquer strategy to mitigate errors and computational costs. Secondly, inspired by the pruning-and-growing111“Pruning” and “Growing” will both be used in following text to refer to rebuild sparsity masks. Pruning changes masks from 1 to 0, whereas growing changes them from 0 to 1. operation [14, 7], we address the limitations of mask selection of pruning in dense networks due to the changeable significance of weights, by re-evaluating weight importance score in sparse networks and rebuilding the sparsity mask through targeted growth of salient weights and pruning of non-salient weights. Thirdly, our analysis reveals that with the advancement of LLMs, mask selection becomes increasingly critical in weight reconstruction. Thus, we prioritize the rebuilding of the sparsity mask and strip the reconstruction of weights. Base on these insights, we propose LLM-Barber, a Block-Aware Rebuilder for Sparsity Mask in One-Shot without any need for fine-tuning or retraining.
@@ III-B Block-Aware Rebuilder for Sparsity Mask
$\mathbf{E}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (4)
[…]
@@
$\mathbf{E}=\mathbf{E_{s}}+\mathbf{\frac{\partial E}{\partial W}}\Delta\mathbf{W}+\frac{1}{2}\Delta\mathbf{W}^{T}\mathbf{H}\Delta\mathbf{W}.$ (5)
The zeroth-order term $\mathbf{E_{s}}$ of the Taylor expansion at sparse networks represents the reconstruction error after mask initialization:
$\mathbf{E_{s}}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M_{i},X)}||_{2}^{2}.$ (6)
@@
where ||$\cdot$|| ⋅ | represents the absolute value operator, and $\mathbf{E}$ denotes the block-aware reconstruction error:
$\mathbf{E}=||\mathbf{\text{Block}(W,X)}-\mathbf{\text{Block}(W\odot M,X)}||_{2}^{2}.$ (9)
@@
Pruning Granularity. Choosing the right pruning granularity is crucial [8]. Traditional methods operate on a layer-wise [15], input-wise [4], or output-wise [8] basis, which mainly address the layer-aware reconstruction problem, known for its output channel independence. Wanda’s output-wise ranking yields superior results compared to other methods. However, in LLM-Barber’s block-aware framework, output channel independence is no longer applicable. Thus, LLM-Barber extends consideration to block-wise granularity, prioritizing all linear layers within a block. Our analysis of the four distinct granularity levels shows that optimal granularity depends on the specific sparse mask initialization, with detailed results discussed in the Ablation Study IV-D.
Mask Rebuilding. With the block-aware reconstruction error, gradient information, sparsity matrix, and granularity established, we proceed to rebuild the sparsity mask for each layer. Consider a cluster of weights $\mathbf{W^{c}}$ under a specific sparsity granularity and its corresponding sparsity mask $\mathbf{M^{c}}$ and pruning metric $\mathbf{S^{c}}$. We define the growing criterion and the pruning criterion:
$gi,gj=\text{argmax}\ \mathbf{S^{c}},\ \text{if}\ \mathbf{M^{c}}_{gi,gj}=0,$ (10)
$pi,pj=\text{argmin}\ \mathbf{S^{c}},\ \text{if}\ \mathbf{M^{c}}_{pi,pj}=1.$ (11)
@@
Our experiments show that LLM-Barber identifies varying proportions of salient weights based on the sparse mask’s initialization method. To regulate mask rebuilding, we introduce a hyperparameter $\alpha$, known as the mask rebuilding ratio. The number of mask rebuilding pairs $N$ is calculated as:
@@
III-C Procedure
Like most post-training pruning methods [15, 8], LLM-Barber is executed within a single global LLM forward pass, with local backward passes for gradient computation in each block. Figure 2 illustrates the LLM-Barber workflow, which consists of four stages:
(1) Sparsity Mask Initialization. Initialize a preliminary sparsity mask from the dense network by a post-training pruning technique. (2) Block-aware Reconstruction Error Computation. We use a block-aware reconstruction error to evaluate the discrepancy between the dense and sparse model outputs. (3) Back-propagation for Gradients. Gradients are automatically derived via back-propagation, and the product of weights and gradients serves as the pruning metric. (4) Sparsity Mask Rebuilding. Masks are sorted by pruning metric, unpruned weights in ascending order while pruned weights in descending order. We rebuild the weight masks by growing newly significant weights and pruning those became non-salient with a specific mask rebuilding ratio.
Output: Rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
[…]
3:for $l$ in $\{1,2,\dots,L\}$ do
4:  $\mathbf{M_{r}}^{l}\leftarrow\mathbf{M_{i}}^{l}$
5:  $\mathbf{S}^{l}\leftarrow|\mathbf{W}^{l}|\cdot|\mathbf{G}^{l}|$
[…]
7:  for $n$ in $\{1,2,\dots,N\}$ do
8:   Obtain growing index $gi,gj$ via Eq.(10)
9:   Obtain pruning index $pi,pj$ via Eq.(11)
10:   $\mathbf{M_{r}}^{l}_{pi,pj}=1$triangleright Weight Growing.
11:   $\mathbf{M_{r}}^{l}_{gi,gj}=0$triangleright Weight Pruning.
12:  end for
13:end for
[…]
15: Identify improvement by comparing $\mathbf{E_{i}}$, $\mathbf{E_{r}}$.
16:return rebuilt sparsity masks $\{\mathbf{M_{r}}^{l}\}^{L}_{l=1}$.
Algorithm 1 Pseudocode of LLM-Barber.
III-D Structured N:M Sparsity
While LLM-Barber primarily targets unstructured sparsity, it can be adapted for structured N:M sparsity. Groups of M weights are pruned to retain only N non-zero weights. During mask rebuilding, it divides each M-group into N pairs (one pruned and one non-pruned weight), then sorts them by output channel to identify mask rebuilding pairs. This method optimizes sparsity mask, leveraging N:M sparsity while maintaining model performance.
TABLE I: WikiText-2 perplexity comparison at 50% sparsity rate. Bold results show improvements of integrating LLM-Barber. Underscored results indicate best performance in each LLM.WR represents weight reconstruction.